Session: Nov 2010 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (roda at Thu Nov 04 15:03:44 2010) Success:

Torun (K. Borkowski at Fri Nov 05 07:20:32 2010) Minor failures: POIN FS
FS hangs at 7:16 UT. Soon recovered but probably no useful observations till 8:05 when frequency dependent pointing offset was adjusted.

Effelsberg (Uwe Bach at Fri Nov 05 11:42:25 2010) Success: ANT
Oil leakage at the elevation gear. Had to stop from 7:40 to 14 UT for repair.

Westerbork (Geert Kuper at Fri Nov 05 19:34:21 2010) Success:
Apart from RFI no known problem.

Onsala (Michael Lindqvist at Sat Nov 06 08:32:05 2010) Success:
RFI, otherwise no known problems.